### v6

# Coprime Bivariate Bicycle Codes and Their Layouts on Cold Atoms

Abstract
Quantum computing is deemed to require error correction at scale to mitigate physical noise by reducing it to lower noise levels while operating on encoded logical qubits. Popular quantum error correction schemes include CSS code, of which surface codes provide regular mappings onto 2D planes suitable for contemporary quantum devices together with known transversal logical gates. Recently, qLDPC codes have been proposed as a means to provide denser encoding with the class of bivariate bicycle (BB) codes promising feasible design for devices.
This work contributes a novel subclass of BB codes suitable for quantum error correction. This subclass employs coprimes and the product $xy$ of the two generating variables $x$ and $y$ to construct polynomials, rather than using $x$ and $y$ separately as in vanilla BB codes. In contrast to vanilla BB codes, where parameters remain unknown prior to code discovery, the rate of the proposed code can be determined beforehand by specifying a factor polynomial as an input to the numerical search algorithm. Using this coprime-BB construction, we found a number of surprisingly short to medium-length codes that were previously unknown. We also propose a layout on cold atom arrays tailored for coprime-BB codes. The proposed layout reduces both move time for short to medium-length codes and the number of moves of atoms to perform syndrome extractions. We consider an error model with global laser noise on cold atoms, and simulations show that our proposed layout achieves significant improvements over prior work across the simulated codes.
1 Introduction
Quantum information is susceptible to errors during storage and operation. As the number of qubits in a quantum circuit increases, so does the frequency of errors. Therefore, quantum error correction (QEC) is a cornerstone of advancing from the current noisy intermediate-scale quantum (NISQ) era to the next era of fault-tolerant quantum (FTQC) computing. Typically, a QEC code is characterized by a 3-tuple $\llbracket n,k,d\rrbracket$, indicating that the code utilizes $n$ physical qubits to encode $k$ logical qubits and can correct up to $\lfloor(d-1)/2\rfloor/$ errors. This encoding process incurs overhead, quantified by the code rate $k/n$. Among the various QEC codes, quantum Low-Density Parity-Check (qLDPC) codes stand out [1, 2, 3, 4, 5] due to their lower-weight stabilizers, low overhead, and high thresholds. As a special case of qLDPC codes, the surface codes, which also feature low-weight stabilizers, are the most commonly used codes in quantum computing due to their simple 2D grid structure and simplicity to perform logical operations [6]. However, surface codes require significant overhead. For example, the rotated surface code has parameters $\llbracket L^{2},1,L\rrbracket$, meaning that it requires $L^{2}$ physical qubits to protect one logical qubit.
In contrast, research has demonstrated the existence of “good” qLDPC codes [3] indicating that qLDPC codes can have $k$ and $d$ scaling linearly as $n$ grows, i.e. codes with parameters $\llbracket n,k=\Theta(n),d=\Theta(n)\rrbracket$. However, having asymptotically good LDPC codes does not necessarily mean having better parameters than codes designed for short to medium lengths. Moreover, the structure of codes can greatly influence the design of hardware and vice versa. Therefore, for practical purposes, people started to seek qLDPC codes that are easy to implement on hardware and have good finite-length performance. For example, Panteleev and Kalachev [7] proposed generalized bicycle (GB) codes along with the BP-OSD decoder that focus on medium-length performance. It is noteworthy that they also discovered the $\llbracket 126,12,10\rrbracket$ GB code, which has the same parameters as the code we will analyze later. Another similar qLDPC code, the BB code [5], has received much attention because they have high thresholds, toric layout, and can be embedded on two planes. Benefiting from the quasi-cyclic and two-thickness properties of BB codes, recent papers have established the feasibility of BB codes on different architectures, including cold atom arrays [8, 9], trapped ion[10], and superconducting [11, 5]. Due to the high connectivity of qLDPC codes, they are considered hard to implement on a superconducting-based platform unless a multi-layer architecture is used [11]. Benefiting from their all-to-all connectivity, cold atom array and trapped ion systems have certain advantages in implementing qLDPC codes. In particular, Viszlai et al. [8] showed that the quasi-cyclic structure of BB codes can be easily implemented using 2D atom array acousto-optic deflectors (AOD) for atom movement.
Constructing BB codes is, nonetheless, a time-consuming process as one has to search for combinations of polynomials to construct a code. Moreover, the parameters of codes can not be guaranteed. This challenge motivated us to develop an algorithm that accelerates the search for good BB codes and even constructs codes with pre-determined parameters. Our work differentiates itself from these approaches by introducing a general and efficient algorithm for searching BB codes in the form described in Eq. (5), which is also the form of codes used by related works. Furthermore, we proposed a novel algorithm that allows us to search for a subclass of BB codes with the desired dimensions, which we call coprime-BB codes. Coprime-BB codes generalize the form in Eq. (5) by allowing mixed terms but also have a restriction of using coprime circulant matrix sizes, making them unattainable through searches limited to the polynomial form of this equation. Thus, coprime-BB codes provide more flexibility for different scenarios on top of the originally proposed BB codes.
To summarize, this work makes the following contributions:
  • We propose a fast numerical algorithm to search for good BB codes by excluding certain polynomial combinations.
  • We propose a new construction of BB codes that allows us to customize the code rate before performing a search, much in contrast to prior search techniques that identified the rate only after returning a new code as a search result. This new method involves selecting two coprime numbers and a factor polynomial, leading us to name this subclass coprime-BB codes.
  • We study the properties of coprime-BB codes, devise a novel method to reduce the search space for unknown codes and develop a new layout tailored for quantum devices utilizing cold atom arrays. The simulations show that the proposed layout achieves a lower logical error rate than under layouts of prior work [8]. This layout benefits from the coprime properties of our codes, which the traditional layout does not exploit, i.e., our coprime codes result in superior performance.
2 Background
2.1 Calderbank-Shor-Steane(CSS) codes
Stabilizer codes are among the most commonly used codes in quantum error correction. One can measure each stabilizer to infer both the type and location of errors in a multi-qubit system. To construct such a code, all stabilizers must commute with each other. Thus, they have a common eigenspace and form a stabilizer group $\mathcal{S}$. The code space defined by such group is
[…]
An $\llbracket n,k,d\rrbracket$ stabilizer code can be defined by nkn-k independent stabilizers, allowing us to encode $k$ qubits of logical information into an $n$-qubit block tolerating up to $\lfloor(d-1)/2\rfloor$} errors. CSS codes are an important class of stabilizer with two sets of stabilizers, $X$-type and $Z$-type, represented by parity-check matrices $H_{X}$ and $H_{Z}$, respectively. Each row in a parity-check matrix corresponds to one stabilizer, and each column corresponds to a qubit. A “1” entry signifies an $X$ or $Z$ operator (depending on whether it is in $H_{X}$ or $H_{Z}$), while a “0” indicates the identity. Consequently, a $X$-type stabilizer acts as $X$ or the identity on each qubit, and a $Z$-type stabilizer acts as $Z$ or the identity on each qubit. Errors can therefore be corrected by handling $Z$ errors and $X$ errors separately. Since all stabilizers must commute with each other, it follows directly that for a CSS code $H_{X}H_{Z}^{T}=0$.
2.2 Bivariate Bicycle Codes
BB codes [5] are a class of CSS codes and LDPC codes. In this context, LDPC means that the stabilizers have bounded weight ensuring low density in both rows and columns of the parity-check matrices.
Let $S_{m}$ be the shift matrix of size $m$, defined as
$S_{m}=I_{m}>>1,$ (2)
where “>>>>” denotes the right cyclic shift for each row, and $I_{m}$ is the $m\times m$ identity matrix. For example,
$S_{3}=\begin{bmatrix}0&1&0\\ 0&0&1\\ 1&0&0\end{bmatrix}$. (3)
By defining $x=S_{l}\otimes I_{m}$ and $y=I_{l}\otimes S_{m}$, it is easy to verify that $xy=yx$ using the mixed-product property of the Kronecker product. This definition forms a bijection from the set of monomials $\{x^{i}y^{j}|0\leq i<l,0\leq j<m\}$ to the set of $(lm)\times(lm)$ matrices generated by $x$ and $y$. Therefore, we can interchangeably use polynomials or monomials in $x$ and $y$ to represent their corresponding matrices. It is also straightforward that $x^{l}=y^{m}=I$. The BB codes can be defined by two polynomials, $A=a(x,y)$ and $B=b(x,y)$, and the parity check matrices for BB code are defined as
$\begin{split}H_{X}&=[A|B]\\ H_{Z}&=[B^{T}|A^{T}].\end{split}$ (4)
and it meets the CSS condition as $H_{X}H_{Z}^{T}=AB+BA=2AB=\bm{0}$ in $\mathbb{F}_{2}$ since $x$ and $y$ commute. In [5], the authors restricted the polynomials to shapes of
$\begin{split}a(x,y)&=x^{a}+y^{b}+y^{c}\\ b(x,y)&=y^{d}+x^{e}+x^{f}.\end{split}$ (5)
Therefore, we can write $A$ and $B$ as $A=A_{1}+A_{2}+A_{3}$ and $B=B_{1}+B_{2}+B_{3}$. Each polynomial has three terms, making each stabilizer supported by six qubits. Besides, we know that $A^{T}=A_{1}^{T}+A_{2}^{T}+A_{3}^{T}=A_{1}^{-1}+A_{2}^{-1}+A_{3}^{-1}$ as $A_{i}$ is the power of $x$ or $y$, which are permutation matrices. Similarly, we have $B^{T}=B_{1}^{-1}+B_{2}^{-1}+B_{3}^{-1}$. It is well known [12] that for any CSS code, $k=n-\text{rank}(H_{X})-\text{rank}(H_{Z})$. For BB codes, according to Lemma 1 in [5], $\text{rank}(H_{X})=\text{rank}(H_{Z})$, so this expression simplifies to
$k=2lm-2\text{rank}(H_{X})=2lm-2\text{rank}(H_{Z}).$ (6)
2.3 Cold Atom Arrays
In cold array-based quantum computers, logical qubits are encoded in a two-dimensional physical atom array, allowing for high-fidelity single- and two-qubit operations to be performed in parallel [13]. Specifically, atoms are loaded into optical traps generated by a spatial light modulator (SLM) and acousto-optic deflectors (AODs). The AOD-based traps enable both vertical and horizontal movement of qubits, thereby providing arbitrary connectivity. A global qubit rotation can be carried out in parallel by illuminating the entire array with Raman excitation. Two-qubit gates can also be executed in parallel by collectively moving qubits. After a global Rydberg laser pulse is applied, entangling gates are performed on pairs of qubits that are brought sufficiently close to each other (closer than the Rydberg blockade radius). The atom array in AOD traps can be moved horizontally, vertically, or even stretched [14], which facilitates the syndrome extraction for BB codes.
In [8], Viszlai et al. propose a layout for BB codes on atom arrays that leverages the quasi-cyclic property of these codes. We will refer to this layout as “BB layout” for the rest of paper; As shown in Eq. (1), one can split the data qubits into two blocks, $L$ and $R$, corresponding to the first $lm$ columns and last $lm$ columns of $H_{X}$ and $H_{Z}$. The ancilla qubits are divided into $X$ and $Z$ blocks corresponding to the rows of $H_{X}$ and $H_{Z}$, respectively. Each qubit can be addressed using a letter from $X,Z,L,R$ with a monomial in $\{x^{i}y^{j}|0\leq i<l,0\leq j<m\}$. This labeling facilitates the process to identify the supports of an ancilla qubit: simply multiply the polynomial associated with the qubit by the relevant monomial by identifying $x^{l}=y^{m}=1$. In other words,
$x^{a}y^{b}\cdot x^{c}y^{d}=x^{(a+c)\text{mod }l}y^{(b+d)\text{mod }m}.$ (7)
For example, given $a(x,y)=1+x$ and $b(x,y)=1+y$, we immediately know that the ancilla qubit $X_{xy}$ supports $L_{xy},L_{x^{2}y},R_{xy}$ and $R_{xy^{2}}$. In addition, we can perform syndrome extraction in parallel by mapping qubits with the same labels, i.e., $X_{x^{a}y^{b}},Z_{x^{a}y^{b}},L_{x^{a}y^{b}},R_{x^{a}y^{b}}$, to a subgrid with two-dimensional coordinates $(a,b)$, as illustrated in Fig. 1. This layout allows any check to be executed by moving ancilla qubits cyclically along the vertical and/or horizontal directions. Qubits near the boundary wrap around to the opposite side, which may require multiple moves to complete the check. For example, if $a(x,y)=x$, we need two steps to complete the syndrome extraction specified by $a(x,y)$ as illustrated in Fig.1 (c-d).
Figure 1: (a) Connections defined by the polynomials $a(x,y)$ and $b(x,y)$ between $X$, $Z$, $L$, and $R$ qubits. (b) The layout from [8] for a BB code with $l=3$ and $m=2$. Arrows indicate the CNOT pairs to be performed when $a(x,y)=x$. In this case, CNOTs are applied between the following $X$ ancilla and $L$ data qubit pairs: $(X_{1},L_{x})$, $(X_{y},L_{xy})$, $(X_{x},L_{x^{2}})$, $(X_{xy},L_{x^{2}y})$, $(X_{x^{2}},L_{1})$, and $(X_{x^{2}y},L_{y})$. (c–d) Lighter orange squares indicate the initial positions of the $X$ checks before movement. Each check monomial requires two moves. (c) Rightward movement enables CNOTs between $(X_{1},L_{x})$, $(X_{y},L_{xy})$, $(X_{x},L_{x^{2}})$, and $(X_{xy},L_{x^{2}y})$. (d) Leftward movement enables CNOTs between $(X_{x^{2}},L_{1})$ and $(X_{x^{2}y},L_{y})$.
3 Proposed Search Algorithms
3.1 Code equivalence
In [5], BB codes are obtained through a numerical search. To accelerate this search, we reduce the search space by eliminating some codes with the same $n,k,d$ parameters.
Equivalence of BB codes —It is straightforward to prove that the following four codes
$\begin{split}\mathcal{C}_{1}:H_{X}&=[A|B],H_{Z}=[B^{T}|A^{T}]\\ \mathcal{C}_{2}:H_{X}&=[A^{T}|B^{T}],H_{Z}=[B|A]\\ \mathcal{C}_{3}:H_{X}&=[B|A],H_{Z}=[A^{T}|B^{T}]\\ \mathcal{C}_{4}:H_{X}&=[B^{T}|A^{T}],H_{Z}=[A|B]\end{split}$ (8)
share the same parameters, allowing us to search within only one class of these codes. These equivalences can also be viewed as special cases of Theorem 6 in [15].
Proof: According to Lemma 1 in [5], every bivariate bicycle code has the same distance over $X$ or $Z$ and $\text{rank}(H_{X})=\text{rank}(H_{Z})$. Let us assume the distances and dimensions of $\mathcal{C}_{1},\dots,\mathcal{C}_{4}$ are $d_{1},\dots,d_{4}$ and $k_{1},\dots,k_{4}$. For CSS codes, $k=n-\text{rank}(H_{X})-\text{rank}(H_{Z})$ and it is easy to see that $k_{1}=k_{4}$ and $k_{2}=k_{3}$ since one can get one code from another by swapping $H_{X}$ and $H_{Z}$. Thus, it is sufficient to prove that
  • $d_{1}=d_{4}$ and $d_{2}=d_{3}$.
  • $k_{1}=k_{2}$ and $d_{1}=d_{2}$.
to ensure $\mathcal{C}_{1},\dots,\mathcal{C}_{4}$ has the same parameters.
Let $\bm{f}$ be the binary string of an arbitrary logical $X$ operator of $\mathcal{C}_{1}$, where $f_{i}=1$ if the operator acts as $X$ on qubit $i$ and $f_{i}=0$ if it acts as $I$. By definition, logical operators commute with every $Z$ stabilizer, thus satisfying
$B^{T}|A^{T}]\bm{f}^{T}=\bm{0}.$ (9)
For the sake of simplicity, given a binary string $\bm{x}$, we use $Z^{\bm{x}}$ to represent a Pauli string $Z^{x_{1}}\otimes Z^{x_{2}}\otimes\cdots\otimes Z^{x_{n}}$. Therefore, Eq. (9) indicates that $Z^{\bm{f}}$ is a logical $Z$ operator of $\mathcal{C}_{4}$ as $B^{T}|A^{T}]$ is the $H_{X}$ of $\mathcal{C}_{4}$. Hence, any $X$ operator of $\mathcal{C}_{1}$ is a $Z$ operator of $\mathcal{C}_{4}$, i.e., $d_{1}\geq d_{4}$. Similarly, $d_{4}\geq d_{1}$. Therefore, $d_{4}=d_{1}$. Using the same reasoning, we can prove that $\mathcal{C}_{2}$ and $\mathcal{C}_{3}$ have the same distance.
Next, we will prove that $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ have the same distance. Let $C_{l}$ be the anti-diagonal matrix of size $l\times l$. We know $C_{l}C_{l}=I$ and $C_{l}MC_{l}=M^{T}$ for any given matrix $M$ of the same size. As $C_{l}$ is a full-rank matrix and
$A^{T}|B^{T}]=C_{lm}[A|B]{\begin{bmatrix}C_{lm}&0\\ 0&C_{lm}\end{bmatrix},}$ (10)
we know $A^{T}|B^{T}]$ and $[A|B]$ have the same rank and $k_{1}=k_{2}$. Let $Z^{\bm{p}}$ be an arbitrary logical $Z$ operator of $C_{2}$, where $\bm{p}=(\bm{\alpha}|\bm{\beta})$ is a length $n$ binary vector and $\bm{\alpha},\bm{\beta}$ are binary vectors of length $n/2$. By definition, we have $A^{T}|B^{T}]\bm{p}^{T}=A\bm{\alpha}^{T}+B\bm{\beta}^{T}=\bm{0}$, i.e.,
$C_{lm}AC_{lm}\bm{\alpha}^{T}+C_{lm}BC_{lm}\bm{\beta}^{T}=\bm{0}.$ (11)
Recall that $C_{lm}C_{lm}=I$. By multiplying both sides with $C_{lm}$, we get
$AC_{lm}\bm{\alpha}^{T}+BC_{lm}\bm{\beta}^{T}=\bm{0},$ (12)
and we know that $(\bm{\alpha}C_{lm}^{T}|\bm{\beta}C_{lm}^{T})$ is a logical $Z$ operator of $\mathcal{C}_{1}$ with the same weight of $\bm{p}$. Therefore, $d_{1}\leq d_{2}$ and, similarly, $d_{2}\leq d_{1}$. Thus, $d_{2}=d_{1}$.∎
We note that these two codes, $\mathcal{C}_{1}:H_{X}=[A|B],H_{Z}=[B^{T}|A^{T}]$ and $\mathcal{C}_{5}:H_{X}=[A^{T}|B],H_{Z}=[B^{T}|A]$, do not always have the same parameters. For example, when $l=6,m=12$, the code constructed by $a(x,y)=x^{4}+y^{2}+y^{6}$ and $b(x,y)=y^{5}+x^{3}+x^{4}$ is a $\llbracket 144,8,10\rrbracket$ code whereas the code constructed by $a(x,y)=x^{2}+y^{6}+y^{10}$ and $b(x,y)=y^{5}+x^{3}+x^{4}$ is a $\llbracket 144,8,8\rrbracket$ code.
3.2 Searching for BB Codes
Based on the equivalence, the accelerated search algorithm is described in Algorithm 1. The function remove_equivalent() removes redundant codes that share the same $\llbracket n,k,d\rrbracket$ parameters as per Eq. (8). The rank() function computes the rank of matrices over $\mathbb{F}_{2}$, and the is_connected() function checks if the code’s Tanner graph is connected. Details of the connectivity test can be found in [5]. We focus on connected Tanner graphs because codes with disconnected Tanner graphs typically have lower distances. This distance_upperbound() function estimates the code distance using a threshold $\tau_{d}$. In essence, the distance can be estimated using any decoding algorithm by applying errors to a code word and checking the decoding result to see if it is a logical error. We conduct multiple trials and track the lowest-weight error that results in a logical error. This gives an upper bound of $d$. Here we used a BP-OSD decoder with 1,000 trials. The threshold $\tau_{d}$ is used to skip the rest of trials whenever an error with weight lower than $\tau_{d}$ is found, saving computation time. We employ a similar approach by introducing a threshold $\tau_{k}$. Since computing the upper bound of $d$ is considerably slower, we first calculate $k$. If $k<\tau_{k}$, the code is discarded without proceeding to the more time-consuming distance calculation. After the search process, we can select codes of interest to calculate their exact distance using an integer programming solver. As mentioned above, BB codes have symmetric distances for $X$ and $Z$ errors. Therefore, it is sufficient to find the $X$ distance to determine the distance of the code. It is important to note that, although our search focused on codes with weight-6 stabilizers of the form specified in Eq.(5), the algorithm can be adapted to search for codes with different weights or forms by modifying its input parameters.
A selection of codes found by Algorithm 1 is shown in Table 1. Notably, Eberhardt et al. [16] propose a $\llbracket 108,16,6\rrbracket$ code and a $\llbracket 162,24,6\rrbracket$ code, which are precisely two and three times the parameters $n,k$ of our $\llbracket 54,8,6\rrbracket$ BB code, respectively. The relationship among these three codes remain an area for future research.
Input : $l,m,\tau_{k},\tau_{d}$
Result: codes of parameters $\llbracket 2lm,\geq\tau_{k},\leq\hat{d}\rrbracket$
Generate all polynomial pairs of form Eq. (5) $L\leftarrow[(a_{1}(x,y),b_{1}(x,y)),$...];
$L^{\prime}\leftarrow\textnormal{{remove\_equivalent}}(L)$ ;
/* Remove codes with the same parameters */
for $i\leftarrow 1$ to $|L^{\prime}|$ do
[…]
    $k\leftarrow 2lm-2\textnormal{{rank}}(H_{X})$;
    if $k<\tau_{k}$ then
       continue ;
    else
       $\hat{d}\leftarrow\textnormal{{distance\_upperbound}}(H_{X},H_{Z},\tau_{d})$;
      end if
 else
    continue ;
   end if
end for
Algorithm 1 Algorithm to search for BB codes
Table 1: Some Novel codes found by Algorithm 1
$l$ $m$ $a(x,y)$ $b(x,y)$ $\llbracket n,k,d\rrbracket$ $kd^{2}/n$
3 9 $1+y^{2}+y^{4}$ $y^{3}+x^{1}+x^{2}$ $\llbracket 54,8,6\rrbracket$ 5.33
7 7 $x^{3}+y^{5}+y^{6}$ $y^{2}+x^{3}+x^{5}$ $\llbracket 98,6,12\rrbracket$ 8.82
3 21 $1+y^{2}+y^{10}$ $y^{3}+x+x^{2}$ $\llbracket 126,8,10\rrbracket$ 6.35
5 15 $1+y^{6}+y^{8}$ $y^{5}+x+x^{4}$ $\llbracket 150,16,8\rrbracket$ 6.83
3 27 $1+y^{10}+y^{14}$ $y^{12}+x+x^{2}$ $\llbracket 162,8,14\rrbracket$ 9.68
6 15 $x^{3}+y+y^{2}$ $y^{6}+x^{4}+x^{5}$ $\llbracket 180,8,16\rrbracket$ 11.38
3.3 Coprime Construction
Based on the commutativity of matrices $x$ and $y$, one can construct valid CSS codes using various polynomial forms other than Eq. (5). For example, we can construct more BB codes by allowing mixed terms or different numbers of pure $x$- and $y$- terms. However, this generalization significantly increases the search space. Especially when using the polynomial form in Eq. (5), finding codes with desirable $k$ and $d$ is already computationally expensive. Thus, we propose a different construction that not only yields new codes and reduces the search space but also guarantees to find codes with a predetermined $k$.
We diverge slightly from the original BB codes by letting $l,m$ be two coprime numbers and $\pi=xy$ to define the polynomials of coprime-BB codes as follows:
$\begin{split}&a(\pi)=\sum a_{i}\pi^{i},\quad b(\pi)=\sum b_{j}\pi^{j},\\ &i,j\in\{0,1,\dots,lm-1\},a_{i},b_{j}\in\{0,1\}.\end{split}$ (13)
Dimension of coprime-BB codes — It is easy to verify that $\langle xy\rangle$ is a cyclic group of order $lm$, thus any monomial in $\{x^{i}y^{j}\>|\>0\leq i<l,\>0\leq j<m\}$ can be expressed as a power of $xy$. Therefore, any polynomials that define the BB code with coprime $l$ and $m$ can be expressed by univariate polynomials $a(\pi)$ and $b(\pi)$ and thus in the form of Eq. (13). Let $g(\pi)=\text{GCD}(a(\pi),b(\pi),\pi^{lm}+1)$, where GCD is the greatest common divisor. Let $\deg g(\pi)$ be the degree of polynomial $g(\pi)$. The BB code defined by $a(\pi)$ and $b(\pi)$ then has dimension
$k=2\deg g(\pi).$ (14)
Proof: As mentioned above, when $l$ and $m$ are two coprime integers, any polynomial in $\mathbb{F}_{2}[x,y]/(x^{l}+1,y^{m}+1)$ can be expressed in $\mathbb{F}_{2}[\pi]/(\pi^{lm}+1)$. We interpret each column of the parity-check matrix as a polynomial by taking the column entries as monomial coefficients. The rest of the proof is similar to Proposition 1 in [7]. Given the column space of $H_{X}$ is equal to
$\begin{split}\text{colsp}(H_{X})&=\{H_{X}\bm{x}|\bm{x}\in\mathbb{F}_{2}^{2lm}\}\\ &=\{A\bm{u}+B\bm{v}|\bm{u},\bm{v}\in\mathbb{F}_{2}^{lm}\},\end{split}$ (15)
it can be represented in terms of polynomials
$\begin{split}\text{colsp}(H_{X})&=\{a(\pi)u(\pi)+b(\pi)v(\pi)|\\ &u(\pi),v(\pi)\in\mathbb{F}_{2}[\pi]/(\pi^{lm}+1)\}.\end{split}$ (16)
Since $R=\mathbb{F}_{2}[\pi]/(\pi^{lm}+1)$ is a univariate polynomial ring, $a(\pi)R$ and $b(\pi)R$ are principal ideals. Thus, $\text{colsp}(H_{X})$ is an principal ideal and is generated by $g(\pi)$ and $\text{rank}(H_{X})=\dim\text{colsp}(H_{X})=lm-\deg g(\pi)$. Therefore, the dimension is given by
$\begin{split}k&=2lm-2\text{rank}(H_{X})\\ &=2lm-2(lm-\deg g(\pi))=2\deg g(\pi).\end{split}$ (17)
The code equivalence described above for BB codes also holds for coprime-BB codes. However, we want to add another rule for coprime-BB codes that is easy to implement and can further reduce the search space by $1/(lm)^{2}$.
Equivalence for coprime-BB code — Let $\mathcal{C}$ be the code defined by polynomials $a(\pi)$ and $b(\pi)$. The code $\mathcal{C}^{\prime}$ defined by the polynomials $\pi^{i}a(\pi)$ and $\pi^{j}b(\pi)$ has the same parameters as $\mathcal{C}$.
Proof: From Eq. (17), we know that $\mathcal{C}$ and $\mathcal{C}^{\prime}$ have the same dimension $k$ because $\pi^{i}$ and $\pi^{j}$ are not factors of $\pi^{lm}+1$, thus multiplying the polynomials by $\pi^{i}$ and $\pi^{j}$ does not affect their greatest common divisor. Assume $\bm{p}=(\bm{\alpha}|\bm{\beta})$ is the binary vector form of a logical $Z$ operator of $\mathcal{C}$, and let $A=a(\pi),B=b(\pi),A^{\prime}=\pi^{i}a(\pi),B^{\prime}=\pi^{j}b(\pi)$ be the matrices corresponding the polynomials. By definition, a logical $Z$ operator satisfies
$A\alpha+B\beta=\bm{0}.$ (18)
Now consider a binary vector $\bm{p}^{\prime}=(\pi^{-i}\bm{\alpha}|\pi^{-j}\bm{\beta})$. Since $\pi$ is a permutation matrix (and, hence, invertible), multiplying by $\pi$ does not change the Hamming weight. Therefore, $\bm{p}^{\prime}$ has the same weight as $\bm{p}$. Moreover, all matrices here commute by construction, implying
$\begin{split}A^{\prime}\pi^{-i}\bm{\alpha}+B^{\prime}\pi^{-j}\bm{\beta}=&A\pi^{i}\pi^{-i}\bm{\alpha}+B\pi^{j}\pi^{-j}\bm{\beta}\\ =&A\alpha+B\beta=\bm{0}.\end{split}$ (19)
Thus, $\bm{p}^{\prime}$ is a logical $Z$ operator of $\mathcal{C}^{\prime}$. Similarly, we can prove this property for a logical $X$ operator by conversely implying that $\mathcal{C}^{\prime}$ and $\mathcal{C}$ are equivalent in terms of their parameters. ∎
Table 2: Some novel coprime-BB codes found by Algorithm 2.
$l$ $m$ $a(\pi),a(x,y)$ $b(\pi),b(x,y)$ $\llbracket n,k,d\rrbracket$ $kd^{2}/n$
3 5 $1+\pi+\pi^{2}=1+xy+(xy)^{2}$ $1+\pi^{2}+\pi^{7}=1+(xy)^{2}+xy^{2}$ $\llbracket 30,4,6\rrbracket$ 4.8
3 7 $1+\pi^{2}+\pi^{3}=1+(xy)^{2}+y^{3}$ $1+\pi^{2}+\pi^{10}=1+(xy)^{2}+xy^{3}$ $\llbracket 42,6,6\rrbracket$ 5.14
5 7 $1+\pi+\pi^{5}=1+(xy)+y^{5}$ $1+\pi+\pi^{12}=1+xy+x^{2}y^{5}$ $\llbracket 70,6,8\rrbracket$ 5.49
2 27 $1+\pi^{3}+\pi^{42}=1+xy^{3}+y^{15}$ $1+\pi^{6}+\pi^{39}=1+y^{6}+xy^{12}$ $\llbracket 108,12,6\rrbracket$ 4
7 9 $1+\pi+\pi^{58}=1+xy+x^{2}y^{4}$ $1+\pi^{13}+\pi^{41}=1+x^{6}y^{4}+x^{6}y^{5}$ $\llbracket 126,12,10\rrbracket$ 9.52
7 11 $1+\pi+\pi^{31}=1+xy+x^{3}y^{9}$ $1+\pi^{19}+\pi^{53}=1+x^{5}y^{8}+x^{4}y^{9}$ $\llbracket 154,6,16\rrbracket$ 9.97
Input : $l,m,\tau_{d},\tau_{k}$
Result: codes of parameters $\llbracket 2lm,\geq\tau_{k},\leq\hat{d}\rrbracket$
$G\leftarrow\textnormal{{factors}}(\pi^{lm}+1)$;
/* Find all factors over $\mathbb{F}_{2}$ */
for $g(\pi)\in G$ do
 if $2\deg g(\pi)<\tau_{k}$ then
    continue;
 $C\leftarrow$ all polynomials $f(\pi)$ in $\mathbb{F}_{2}[\pi]/(\pi^{lm}+1)$ s.t. $\text{wt}(f(\pi))=3$;
 /* Or > 3 to get higher weight codes */
 $C^{\prime}\leftarrow$ all polynomials $c(\pi)$ in $C$ s.t. $c(\pi)\mod g(\pi)=0$;
 $L\leftarrow$ all combinations $(a(\pi),b(\pi))$ chosen from $C^{\prime}$ s.t. $\text{GCD}(a(\pi),b(\pi))=g(\pi)$;
 $L^{\prime}\leftarrow\textnormal{{remove\_equivalent}}(L)$;
 for $i\leftarrow 1$ to $|L^{\prime}|$ do
[…]
    $k\leftarrow 2\deg g(\pi)$;
    $\hat{d}\leftarrow\textnormal{{distance\_upperbound}}(H_{X},H_{Z},\tau_{d})$;
    else
       continue ;
      end if
   end for
end for
Algorithm 2 An algorithm to search for BB codes with the new form of polynomials.
Coprime-BB codes can also be viewed as a special case of generalized bicycle (GB) codes [7]. In GB codes, polynomials are identified as sums of cyclic shift matrices. In contrast, coprime-BB codes identify polynomials as sums of the Kronecker products of two cyclic shift matrices with coprime dimensions. By adapting the new construction to Algorithm 1, we propose Algorithm 2. The latter algorithm results in a significantly reduced search space as only coprimes and codes with desired $k$ are being considered. Using Algorithm 2, we found a number of interesting coprime-BB codes, which are shown in Table 2 with their polynomials in the bivariate form. We also noted that in [7], the author found a $\llbracket 126,12,10\rrbracket$ with different polynomials, i.e., we are not the first to discover this particular code, yet our search algorithm identified it based on a different basis. For large $l,m$, the results are not exhaustive as numerous polynomials meet the condition $\text{GCD}(a(\pi),b(\pi))=p(\pi)$. As the algorithm outputs codes with the same $k$, the results shown were obtained by running the search algorithm for a short duration and selecting the output code with the highest estimated $d$.
Additionally, we visualize our newly discovered codes alongside those proposed in [5] in Fig. 2 (restricted to $n<200$). The codes are plotted using the metric $kd^{2}/n$ against the code length $n$, which highlights improvements over surface codes, characterized by $n\propto kd^{2}$. As shown in the figure, our new constructions expand the spectrum of available codes at short lengths, offering greater flexibility in code selection suitable for near-term module sizes suitable for distributed quantum computing.
We observe that the $\llbracket 126,12,10\rrbracket$ code offers the highest $d$ and one of the highest $k$, which is comparable to the $\llbracket 144,12,12\rrbracket$ “gross” code proposed in [5], making it suitable for scenarios where error rates are moderate to high, and ample qubit resources exist. The $\llbracket 42,6,6\rrbracket$ code has a low $d$ but provides the highest rate of these codes and is best suited for scenarios with low error rates or where hardware resources are limited. The $\llbracket 70,6,8\rrbracket$ code offers a balanced trade-off between error correction capabilities and code length, making it a versatile option for environments with moderate physical error rates and resource constraints.
Figure 2: Comparison of quantum codes using the normalized metric $kd^{2}/n$ versus code length $n$ for $n<200$. The plot includes codes from [5] (“Codes in [5]”), our constructions matching the form in [5], and newly discovered coprime-based codes.
4 A Novel Layout Optimization for Coprime-BB Codes
As a subclass of BB codes, coprime-BB codes distinguish themselves by one major feature: Almost every monomial in the polynomials that define them is “mixed”, i.e., each monomial contains both $x$ and $y$. In the BB layout by [8], we need four steps to perform the CNOTs required by a mixed monomial as shown in Fig. 3. This is because the mixed term check introduces two types of periodicity, one vertically and one horizontally, rather than just one type of periodicity for codes defined by Eq. (5). However, a global Rydberg laser pulse introduces errors to all atoms after each step of movement, not just those within the blockade radius [14]. Consequently, codes with mixed terms are more prone to error, since checking them requires additional global laser pulses and thereby increasing the likelihood of errors on idle qubits.
Figure 3: Movements required to perform an xyxy check ($l=3,m=2$) on the BB layout from [8]. The required CNOTs are between the following pairs: $(X_{1},L_{xy})$, $(X_{y},L_{x})$, $(X_{x},L_{x^{2}y})$, $(X_{xy},L_{x^{2}})$, $(X_{x^{2}},L_{y})$, and $(X_{x^{2}y},L_{1})$. Four movement steps are needed, as two periodic boundary conditions are crossed during the process.
Despite this drawback, coprime-BB codes have another useful property: Every monomial is of the form of $(xy)^{i}$. This structure naturally suggests a one-dimensional arrangement of atoms, which can avoid multiple movements and thereby help reduce the error rates. Therefore, we propose a novel layout design for coprime-BB codes using this property. As illustrated in Fig. 4, we organize the $X,Z,L$ and $R$ qubits with the same subscript vertically, and we place them horizontally in the order of subscripts $1,xy,(xy)^{2},...,(xy)^{lm-1}$. By doing so, every move of form $(xy)^{i}$ can be seen as a cyclic shift horizontally, i.e., it can be performed in two moves: Move left by lmilm-i subgrids and then move right by $lm$ subgrids. We call this novel layout the “CBB layout” for the remainder of this paper. Since $l$ and $m$ are coprime integers, we can ensure that this layout covers all qubits without repetition in $lm$ columns. As a result, any syndrome extraction can be accomplished by moving the ancilla qubits horizontally (if we ignore the minor vertical movement between the atoms with the same subscript).
Figure 4: The movements to perform a xyxy check using the proposed CBB layout ($l=3,m=2$).
It is also worth noting that the move distance to perform a mixed-term check in the BB layout is at least $4(l+m)$, assuming that the atom distance is 1. In contrast, the move distance of the CBB layout is at least $2lm$. Notice that the latter term starts lower but grows faster than the former as the code length increases, but the move-time cross-over point is beyond the scale of studied code sizes today, i.e., in practice CBB layouts perform better for realistic codes). Figure 5 compares the corresponding movement times for coprime-BB codes of different layouts, where a cycle refers to the time needed for one syndrome-extraction cycle, and a round consists of $d$ such cycles. We also add the movement time of the $\llbracket 144,12,12\rrbracket$ code under scheduling proposed in [8] as a reference. Following the approach in [17, 8], we calculate movement times assuming an atom spacing of $5\mu m$ and an acceleration of $0.02\mu m/\mu s^{2}$. The time for a $(\Delta x,\Delta y)$ move, which proceeds along along Manhattan path, is given by
$\sqrt{6\Delta x/0.02}+\sqrt{6\Delta y/0.02}.$ (20)
The optimal route to complete all monomial checks is computed using a traveling salesman problem (TSP) solver on the Manhattan distance metric. From Figure 5 we observe that the proposed CBB layout has a lower cycle time despite having a higher move distance. This is because the BB layout requires many short moves, whereas the CBB layout relies on fewer but longer moves, which is more efficient when factoring in acceleration overhead, i.e., AOD movement first accelerates to the mid-point before it slows down as suggested in Eq. (20). This benefit of CBB is becoming smaller as the code length becomes large, but we have not observed this phenomenon in the codes tested here.
Discussion: In the CBB layout, qubits are arranged into long stripes forming a $1\times lm$ grid, rather than a $l\times m$ rectangular array. This geometry places increased demands on acousto-optic deflectors, since the larger number of columns requires a correspondingly higher number of resolvable radio-frequency tones, which may constrain the maximum achievable array size due to finite AOD bandwidth and spot-resolution limits. However, coprime-BB codes are known to be asymptotically limited [18], and are therefore not intended for very large code distances. As a result, our focus is on short-to-medium code lengths, for which the required number of sites per row remains moderate. Recent experimental demonstrations of neutral-atom arrays have achieved grids as large as $120\times 12$ [19] suggesting that resolving on the order of $10^{2}$ atoms per row is feasible with current technology.
In addition, ancilla transport in the CBB layout typically involves longer distances per move than in BB layouts. Although our analysis indicates that the total number of movement operations, as well as the overall transport time for the codes considered here, is reduced, this reduction is achieved at the cost of higher average transport velocities. As discussed in [20, 21], the start-stop jerking effect could be a major source of heating and atom loss. The trade-off between fewer start–stop events and increased velocity makes the net impact on transport-induced noise, such as motional heating, difficult to assess without detailed experimental characterization. In summary, the CBB layout is more suitable for short-to-medium code lengths on near-term devices.
Figure 5: Movement time of different codes and layouts. The CBB layout for $\llbracket 144,12,12\rrbracket$ code is omitted because the CBB layout only applies for coprime-BB codes.
5 Numerical Results
In this section, we evaluate the error rates of the newly found codes through theoretical analysis and numerical simulation. We begin by illustrating the performance of the codes under the code-capacity model using Monte Carlo simulations, followed by a simulation based on the circuit-based noise model using Stim [22]. All simulation results are gathered for 100 or more logical errors, ensuring that error bars remain below approximately $\pm 10\%$.
5.1 Code-Capacity Model
In the code-capacity model, we assume that all gate operations and measurements are perfect. $X,Y$ and $Z$ errors are applied with $p/3$ independently for each data qubit. In Fig. 6, the logical error rates (y-axis) of proposed coprime-BB codes and the $\llbracket 144,12,12\rrbracket$ BB code are depicted for different physical error rates (x-axis). The decoder used in the simulation is BP-OSD [7, 23] with a maximum of 10,000 min-sum (MS) iterations, variable scaling factor, and the “OSD_CS” method of order 10. As expected, codes with a larger distance $d$ generally exhibit better performance than those with a smaller $d$. Among codes with the same distance $d$, such as $\llbracket 108,12,6\rrbracket$, and $\llbracket 30,4,6\rrbracket$, the code with shorter code length tends to have a lower logical error rate as longer codes have more sources of errors.
Figure 6: Logical error rates of different codes under code capacity model. The lowest error rate point for $\llbracket 144,12,12\rrbracket$ is simulated under physical error rate $p=1.5\times 10^{-2}$.
5.2 Circuit-Level Noise Model On Cold Atoms Arrays
Under the circuit-level noise model, we consider a more realistic scenario in which errors can occur during any operation, except for classical processing. We adopt the common practice as described in [8], where the syndrome extraction is performed by multiple rounds to compensate for measurement errors. In each round, noise is introduced independently on each measurement and gate operation with a certain probability. The number of rounds is set equal to the distance of the code. After performing the desired rounds, the syndrome history is fed into the decoder to estimate the final error. In our simulations, we assume that the single-qubit gate error rate ($e_{1g})$, two-qubit gate error rate ($e_{2g})$, and readout error rate ($e_{r}$) all share a common probability $p$. We model single- and two-qubit gate errors using depolarizing channels: for single-qubit gates, each non-identity Pauli error occurs with probability $p/3$, while for two-qubit gates, it occurs with probability $p/15$. Idle errors are applied after each move and are determined based on the device’s relaxation time ($T_{1}$), dephasing time ($T_{2}$), and the time required for each move (see [8] for details). We will use $T_{1}=T_{2}=1\>s$ throughout this work. Read-out errors are applied using $X$ or $Z$ flip with probability $p$ based on the measurement basis before each measurement.
According to [14], a global Rydberg laser introduces noise on all atoms, including those not actively participating in entanglement operations. To model this effect, we introduce a coefficient $c$. Each time a two-qubit global gate is applied, in addition to the two-qubit depolarization noise on the interacting qubit pairs, we apply an additional single-qubit depolarization noise with probability $c\cdot e_{2g}$ to all qubits.
Figure 7: Logical error rates per round ($p_{L}$) of selected codes under circuit-level noise model with different physical error rate ($p$).
Figure 8: The circuit level simulations for different coprime-BB codes under different coefficients $c$ and using different layouts. $p_{L}$ is logical error rate per round and $p$ is physical error rate. (a-f) share the same legend as in (a). (a-c) show the comparison for lower distance coprime-BB codes, and (d-f) show that for codes with higher distances.
Fig. 7 presents the logical error rate for selected coprime-BB codes across a range of physical qubit error rates. Except for the $\llbracket 144,12,12\rrbracket$ BB code, which is incompatible with the CBB layout, all codes are arranged in the CBB layout. A global layer error rate coefficient of $c=0.1$ is used for all codes. Unlike the code-capacity model, we evaluate performance via the logical error rate per cycle as
$p_{L}=1-(1-\frac{N_{e}}{N})^{1/d},$ (21)
where $N_{e}$ is the number of logical errors and $N$ is the number of simulations.
Similar to the code-capacity model, codes with higher $d$ tend to achieve a lower logical error rate per cycle. However, we observe that the $\llbracket 42,6,6\rrbracket$ code exhibits a lower logical error rate than $\llbracket 30,4,6\rrbracket$, which has a shorter code length and the same $d$. This discrepancy arises because they have different circuit distances $d_{circ}$, defined as the minimum number of error mechanisms to flip a logical observable without triggering any detection event. Since errors can propagate through the circuit and effectively increase the number of errors, we usually have $d_{circ}<d$. Specifically, $\llbracket 30,4,6\rrbracket$ has $d_{circ}=3$, whereas $\llbracket 42,6,6\rrbracket$ has $d_{circ}=4$. In Fig. 7, we also observe that the performance gap between codes $\llbracket 144,12,12\rrbracket$ and $\llbracket 126,12,10\rrbracket$ is much smaller than in Fig. 6. This reduced gap is likely due to the same reason, i.e., simulations indicate that the $\llbracket 144,12,12\rrbracket$ has a circuit-level distance of $d_{circ}\leq 10$, while the $\llbracket 126,12,10\rrbracket$ has $d_{circ}\leq 9$.
Figure 8 presents the logical error rates $p_{L}$ for different coprime-BB codes under both the BB and CBB layouts for global laser error coefficients $c\in\{0.1,0.2,0.5\}$. Across all tested codes, the proposed CBB layout consistently achieves lower logical error rates than the BB layout, owing primarily to its shorter round/cycle time and fewer ancilla movements. These advantages reduce the number of required global CNOT operations, thereby lowering the accumulated noise level on idle qubits. The improvement is more significant for high distance codes, especially when the global noise coefficient $c$ is high. For instance, we have an error rate reduction around $1/10$ on the $\llbracket 126,12,10\rrbracket$ and $\llbracket 154,6,16\rrbracket$ codes for $c=0.5$. In contrast, at $c=0.1$, the improvement is $1/2$ and $1/6$, respectively. This is due to a higher $c$ indicating that each global laser pulse introduces more noise on all qubits. As the CBB layout saves two global laser operations per monomial in a polynomial, we can expect even greater gains for codes with higher polynomial weight.
6 Related Work
Recently, various constructions based on qLDPC codes have been proposed to achieve different objectives. Koukoulekidis et al. [24] proposed algebraic extensions to expand a small GB code into a family of larger GB codes by selecting a sequence of expansion factors. The authors also introduced scalable codes that embed the original short codes into extended codes, enhancing scalability in superconducting architectures. Voss et al. [25] expanded the concept of BB codes by introducing an additional type of indeterminate variable, leading to the creation of trivariate bicycle (TB) codes. These new codes reduce the weight of stabilizers from 6 to 4-5, making them more practical for hardware implementation. However, it is important to note that some of these codes exhibit a lower rate or distance compared to BB codes of similar length. Eberhardt et al. [16] investigated the algebraic structure of BB codes and uncovered certain symmetry properties. These properties allow for the explicit construction of logical operators and certain fault-tolerant gates for BB codes. Shaw et al. [26] proposed a “morphing” circuit design for syndrome extraction on BB codes. The proposed circuit design has only six rounds of CNOT gates instead of seven [5]. By applying the proposed circuit, the authors discovered a new family of BB codes, including codes that have the same $\llbracket n,k,d\rrbracket$ parameters as [5]. The authors also provide a sufficient condition for the circuit to be applied to the other two-block group algebra (2BGA) [15] codes.
7 Conclusion and Future Work
We developed fast numerical search algorithms to discover BB codes and introduced a novel construction method using factor polynomials of $\mathbb{F}_{2}[\pi]/(\pi^{lm}+1)$, where $l$ and $m$ are coprime integers. The new construction enables us to know the rate of BB codes before constructing them. We also introduced a new error model that accounts for noise of global Rydberg laser pulses affecting non-interacting qubits and demonstrated the error rates of our newly discovered codes in simulation. Moreover, we devise a novel and more efficient mapping of coprime-BB codes tailored to cold atom-array architectures. Our approach achieves shorter movement times, fewer moves, and lower error rates on the codes we tested. These features make the coprime-BB code a candidate for quantum memory. However, further research is needed to explore other properties of these codes, such as logical gate constructions and implementation on superconducting architectures.
Acknowledgment
The authors would like to thank Joshua Viszlai, Hanrui Wang, and John Stack for fruitful discussions on cold atom arrays and error correction. They also thank Yu-An Chen for identifying errors in the manuscript. This work was supported in part by NSF awards MPS-2410675, PHY-1818914, PHY-2325080, MPS-2120757, CISE-2217020, and CISE-2316201 as well as DOE DE-SC0025384.
Appendix A Other BB Codes Found
In addition to the BB codes provided so far, we present additional coprime-BB codes found by Algorithm 1. Any “obviously” inferior codes found are not included in this table, e.g., codes with the same $n$ but lower $k$ or $d$.
Table 3: Selected weight-6 codes found by Algorithm 1.
$l$ $m$ $a(x,y)$ $b(x,y)$ $\llbracket n,k,d\rrbracket$ $kd^{2}/n$
3 3 $1+x+y$ $1+x^{2}+y^{2}$ $\llbracket 18,4,4\rrbracket$ 3.56
3 6 $1+y+y^{2}$ $x^{3}+y+y^{2}$ $\llbracket 36,8,4\rrbracket$ 3.56
3 6 $x+y^{2}+y^{3}$ $1+y+x^{2}$ $\llbracket 36,4,6\rrbracket$ 4
3 9 $x+y+y^{3}$ $1+y^{2}+x^{2}$ $\llbracket 54,4,8\rrbracket$ 4.74
7 14 $1+y+y^{3}$ $y^{7}+x+x^{3}$ $\llbracket 196,18,8\rrbracket$ 5.87
Appendix B Other Coprime-BB Codes Found
Table 4: Selected weight-6 codes found by Algorithm 2
$l$ $m$ $a(\pi),a(a,y)$ $b(\pi),b(x,y)$ $\llbracket n,k,d\rrbracket$ $kd^{2}/n$
2 7 $1+\pi^{1}+\pi^{3}=1+xy+xy^{3}$ $1+\pi^{1}+\pi^{10}=1+xy+y^{3}$ $\llbracket 28,6,4\rrbracket$ 3.43
2 9 $1+\pi^{2}+\pi^{10}=1+y^{2}+y$ $1+\pi^{4}+\pi^{8}=1+y^{4}+y^{8}$ $\llbracket 36,8,4\rrbracket$ 3.56
3 7 $1+\pi+\pi^{5}=1+xy+x^{2}y^{5}$ $1+\pi^{2}+\pi^{10}=1+(xy)^{2}+xy^{3}$ $\llbracket 42,10,4\rrbracket$ 3.8
3 8 $1+\pi+\pi^{2}=1+xy+(xy)^{2}$ $1+\pi^{2}+\pi^{10}=1+(xy)^{2}+xy^{2}$ $\llbracket 48,4,8\rrbracket$ 5.33
3 10 $1+\pi^{2}+\pi^{8}=1+(xy)^{2}+x^{2}y^{8}$ $1+\pi^{4}+\pi^{16}=1+xy^{4}+xy^{6}$ $\llbracket 60,16,4\rrbracket$ 4.27
3 11 $1+\pi+\pi^{5}=1+xy+x^{2}y^{5}$ $1+\pi+\pi^{23}=1+xy+x^{2}y$ $\llbracket 66,4,10\rrbracket$ 6.06
4 7 $1+\pi+\pi^{3}=1+xy+(xy)^{3}$ $1+\pi^{5}+\pi^{11}=1+xy^{5}+x^{3}y^{4}$ $\llbracket 56,6,8\rrbracket$ 6.86
5 9 $1+\pi+\pi^{4}=1+xy+(xy)^{4}$ $1+\pi^{8}+\pi^{34}=1+x^{3}y^{8}+x^{4}y^{7}$ $\llbracket 90,4,12\rrbracket$ 6.4
5 9 $1+\pi+\pi^{12}=1+xy+x^{2}y^{3}$ $1+\pi^{2}+\pi^{9}=1+(xy)^{2}+x^{4}$ $\llbracket 90,8,8\rrbracket$ 5.69
6 7 $1+\pi+\pi^{3}=1+xy+(xy)^{3}$ $1+\pi^{8}+\pi^{31}=1+x^{2}y+xy^{3}$ $\llbracket 84,6,10\rrbracket$ 7.14
6 11 $1+\pi+\pi^{2}=1+xy+(xy)^{2}$ $1+\pi^{11}+\pi^{28}=1+x^{5}+x^{4}y^{6}$ $\llbracket 132,4,14\rrbracket$ 5.94
[…]
7 9 $1+\pi^{4}+\pi^{19}=1+(xy)^{4}+x^{5}y$ $1+\pi^{6}+\pi^{16}=1+(xy)^{6}+x^{2}y^{7}$ $\llbracket 126,6,14\rrbracket$ 9.33
9 10 $1+\pi+\pi^{4}=1+xy+(xy)^{4}$ $1+\pi^{23}+\pi^{62}=1+x^{5}y^{3}+x^{8}y^{2}$ $\llbracket 180,8,16\rrbracket$ 11.38
Table 5: Selected weight-8 codes found by Algorithm 2
$l$ $m$ $a(\pi),a(x,y)$ $b(\pi),b(x,y)$ $\llbracket n,k,d\rrbracket$ $kd^{2}/n$
3 4 $1+\pi+\pi^{3}+\pi^{4}=1+xy+y^{3}+x$ $1+\pi^{2}+\pi^{5}+\pi^{9}=1+(xy)^{2}+x^{2}y+y$ $\llbracket 24,8,4\rrbracket$ 5.33
3 5 $1+\pi+\pi^{2}+\pi^{7}=1+xy+(xy)^{2}+xy^{2}$ $1+\pi+\pi^{4}+\pi^{10}=1+xy+xy^{4}+x$ $\llbracket 30,10,4\rrbracket$ 5.33
3 5 $1+\pi+\pi^{3}+\pi^{4}=1+xy+y^{3}+xy^{4}$ $1+\pi+\pi^{3}+\pi^{7}=1+xy+y^{3}+xy^{2}$ $\llbracket 30,6,5\rrbracket$ 5
3 7 $1+\pi+\pi^{3}+\pi^{13}=1+xy+y^{3}+xy^{6}$ $1+\pi+\pi^{4}+\pi^{9}=1+xy+xy^{4}+y^{2}$ $\llbracket 42,12,5\rrbracket$ 7.14
3 7 $1+\pi+\pi^{3}+\pi^{4}=1+xy+y^{3}+xy^{4}$ $1+\pi+\pi^{6}+\pi^{10}=1+xy+y^{6}+xy^{3}$ $\llbracket 42,6,7\rrbracket$ 7
3 8 $1+\pi+\pi^{2}+\pi^{3}=1+xy+(xy)^{2}+y^{3}$ $1+\pi^{3}+\pi^{9}+\pi^{14}=1+y^{3}+y+x^{2}y^{6}$ $\llbracket 48,6,8\rrbracket$ 8
3 8 $1+\pi+\pi^{3}+\pi^{10}=1+xy+y^{3}+xy^{2}$ $1+\pi^{3}+\pi^{7}+\pi^{16}=1+y^{3}+xy^{7}+x$ $\llbracket 48,10,6\rrbracket$ 7.5
4 5 $1+\pi+\pi^{6}+\pi^{15}=1+xy+x^{2}y+x^{3}$ $1+\pi^{2}+\pi^{5}+\pi^{7}=1+(xy)^{2}+x+x^{3}y^{2}$ $\llbracket 40,14,4\rrbracket$ 5.6
4 5 $1+\pi+\pi^{2}+\pi^{3}=1+xy+(xy)^{2}+(xy)^{3}$ $1+\pi+\pi^{3}+\pi^{10}=1+xy+(xy)^{3}+x^{2}$ $\llbracket 40,6,6\rrbracket$ 5.4
4 5 $1+\pi+\pi^{4}+\pi^{5}=1+xy+y^{4}+x$ $1+\pi+\pi^{4}+\pi^{9}=1+xy+y^{4}+xy^{4}$ $\llbracket 40,8,5\rrbracket$ 5
4 7 $1+\pi+\pi^{2}+\pi^{4}=1+xy+(xy)^{2}+y^{4}$ $1+\pi^{2}+\pi^{6}+\pi^{19}=1+(xy)^{2}+x^{2}y^{6}+x^{3}y^{5}$ $\llbracket 56,8,8\rrbracket$ 9.14
4 7 $1+\pi+\pi^{4}+\pi^{9}=1+(xy)+y^{4}+xy^{2}$ $1+\pi+\pi^{17}+\pi^{20}=1+xy+xy^{3}+y^{6}$ $\llbracket 56,14,6\rrbracket$ 9
5 6 $1+\pi+\pi^{2}+\pi^{7}=1+xy+(xy)^{2}+x^{2}y$ $1+\pi^{3}+\pi^{12}+\pi^{25}=1+(xy)^{3}+x^{2}+y$ $\llbracket 60,12,7\rrbracket$ 9.80
5 6 $1+\pi+\pi^{3}+\pi^{4}=1+xy+(xy)^{3}+(xy)^{4}$ $1+\pi^{2}+\pi^{11}+\pi^{18}=1+(xy)^{2}+xy^{5}+x^{3}$ $\llbracket 60,6,9\rrbracket$ 8.1
5 7 $1+\pi+\pi^{2}+\pi^{4}=1+xy+(xy)^{2}+(xy)^{4}$ $1+\pi+\pi^{6}+\pi^{24}=1+xy+xy^{6}+x^{4}y^{3}$ $\llbracket 70,8,9\rrbracket$ 9.26
In addition to the coprime-BB codes provided so far, we present other coprime-BB codes found by Algorithm 2 in Table 4 and 5.